OpenStudy (anonymous):

How is it that (2^n)/(2^n-1) simplifies to : (1)/(1-(1/2)^n)

OpenStudy (anonymous):

easier to work backwards probably to see it

OpenStudy (anonymous):

\[\frac{1}{1-\frac{1}{2^n}}\] \[\frac{1}{\frac{2^n-1}{2^n}}\] \[\frac{2^n}{2^n-1}\]

OpenStudy (turingtest):

divide top and bottom by 2^n

OpenStudy (anonymous):

thank you I got it now.